OpenStudy (anonymous):

A cyclist rode his bike for 4 hours. In that time, he rode 70 miles. What was his speed in miles per hour? 0.057 miles per hour 17.5 miles per hour 28 miles per hour

OpenStudy (whpalmer4):

70 miles divided by 4 hours. How many miles traveled in each hour, assuming that the speed is constant?

OpenStudy (anonymous):

17.5?

OpenStudy (whpalmer4):

Let's check: if you ride for 4 hours at 17.5 miles/hour, how far do you go? \[d = rt\]\[d = 17.5\frac{\text{ miles}}{\text{ hour}}*4\text{ hours} = 17.5*4 \frac{\text{ miles}}{\cancel{\text{ hour}}}\cancel{\text{ hours}} = 70 \text{ miles}\]Looks good!

OpenStudy (anonymous):

thanks :)

OpenStudy (whpalmer4):

You can also quickly see that neither of the other two answers is reasonable. If we only go 0.057 miles in an hour, we have to travel more than 16 hours just to go the first mile! In fact, that's only about an inch each second, or about 9 times the speed of the fastest garden snail. If we try 28 miles per hour, 4*28 is about the same as 4*30, so we're looking at well over 100 miles traveled. Clearly not the right answer, either.